A company wants to mix peanuts which contain 15% protein and cashews, which contain 40% protein to make a trail mix. Suppose the company wants to make 1000 pounds of trail mix that contain 27% protein. How many pounds of cashew must be used?
Answer by mananth(16949)

A company wants to mix peanuts which contain 15% protein and cashews, which contain 40% protein to make a trail mix. Suppose the company wants to make 1000 pounds of trail mix that contain 27% protein. How many pounds of cashew must be used?
...
peanuts --------------- 15% ---------- x pounds
cashews-----------------40%------------1000-x pounds
trail mix---------------27%------------1000
...
Mix equation
15x+40(1000-x)=27*1000
15x+40,000-40x=27,000
-25x=-13,000
x=-13,000-25
x=520 pounds of peanuts
1000-520=480 pounds of cashews
..
m.ananth@hotmail.ca